multiply this problem
(2a+2)^3=

Answer by jim_thompson5910(35256)

(2a+2)^3 = (2a+2)(2a+2)(2a+2)

(2a+2)^3 = (2a+2)(4a^2 + 8a + 4)

(2a+2)^3 = 2a(4a^2 + 8a + 4) + 2(4a^2 + 8a + 4)

(2a+2)^3 = 8a^3 + 16a^2 + 8a + 8a^2 + 16a + 8

(2a+2)^3 = 8a^3 + 24a^2 + 24a + 8